OpenStudy (anonymous):

Alex and his sister Gina were asked to mow the lawn. Each of them are to mow half of the lawn. The lawn is rectangular in shape and measure 120'x58'. If Alex started in one corner and went around the edges of the lawn how many times would he have to mow around the edges to mow to his half.

OpenStudy (anonymous):

i dont think its stated well

OpenStudy (amistre64):

depends on the width of the mower

OpenStudy (anonymous):

the last sentence, makes no sense

OpenStudy (anonymous):

no but the way I am thinking about it is they have a big rectangle representing the whole thing , then you need to calculate the width from the edge he must mow, to split it into two equal sections

OpenStudy (radar):

Let the mower width be x.

OpenStudy (anonymous):

oh Im so sorry the mower cuts a 2foot wide path

OpenStudy (anonymous):

ohh well that makes it bout 1000times easier :P

OpenStudy (amistre64):

a 24inch cut eh .. is it a john deere ?

OpenStudy (amistre64):

do we make a perfect cut so that each pass is exactly on the cut line of the one before? or do we have to 1/4 step it over to make sure we get it all?

OpenStudy (amistre64):

how sharp is the blade, cause i always get peices left over that defy all cutting conventions :)

OpenStudy (anonymous):

@amistre64; i dont know the question did not say

OpenStudy (radar):

There is 6960 square feet for the entire lawn so Alex has 1/2 of that to mow or 3480 sq ft. First orbit would be 356x square feet. If the mower is 9.8 feet wide that would do it lol.

OpenStudy (radar):

Oh, it is 2 ft. wide That makes the problem interesting as each orbit is smaller. So much for the 9 ft mower!

OpenStudy (amistre64):

3480 is the area we wanna cut; each pass on a perimeter takes off l*2' 2(120-0)+ 2(58-2) + 2(120-2) + 2(58-4) + 2(120-4) + 2(58-6) + 2(120-6) + 2(58-8) ..etc

OpenStudy (amistre64):

2(178) - 2(1) + 2(178) - 2(3) + 2(178) - 2(5) + 2(178) - 2(7) ......

OpenStudy (amistre64):

356n - 2(1+3+5+7+9+11+13 ...)

OpenStudy (amistre64):

B{n} = 356n + 2(2n -1) ... right?

OpenStudy (amistre64):

+2 means -2 :)

OpenStudy (anonymous):

@mrain do u know the answer ? my answer is way to big.......

OpenStudy (anonymous):

is it something abt 387

OpenStudy (radar):

Looks like he would do more than his half on the completion of his 6th orbit.

OpenStudy (anonymous):

myne answer for the orbit was abt something 9.87

OpenStudy (anonymous):

i do not know the answer, very sorry, wish i did

OpenStudy (anonymous):

whats the area for the first orbit

OpenStudy (anonymous):

@radar; thats what i got

OpenStudy (anonymous):

myne was 384= 120*2 + (58-4)*2

OpenStudy (anonymous):

it would take his about 5.5 orbits to equal almost half but it wont equal the exact half, thats what i am understanding

OpenStudy (anonymous):

i used AP........ 384 is the first term and every time area gets reduced by 16 then by formula .... 3840 = 384 +(n-1)(-16)...where n should be number of trip.....pls do tell me if i am wrong....

OpenStudy (anonymous):

sounds right but i truly dont know what is actually right

OpenStudy (anonymous):

i dont think that works because you come out to a negative

OpenStudy (amistre64):

352 + (352-2n)(n-1) 352-2n n-1 -------- 352n -2n^2 -352 +2n ------------ 354n -2n^2 that should do it :) 2380 = 354n - 2n^2 2n^2 -354n +2380 = 0 354/4 +- sqrt(354^2 -4(2380)(2))/4 354/4 +- sqrt(106276)/4 354 + 326 ---------- = 680/4 = 170 4 354 - 326 ---------- = 28/4 = 7 ... id say 7 passes 4

OpenStudy (amistre64):

23 352 +348 +344 +340 +336 +332 +328 ----- 2380 ; 7 passes

OpenStudy (anonymous):

whats 352?

OpenStudy (amistre64):

2(120-0+58-2) = 352 is the area of going half way around

OpenStudy (amistre64):

2(120-2+58-4) = 2(120+58-6)=344 .... bummer forgot how to add :)

OpenStudy (anonymous):

dont u think it should be -4 not -2

OpenStudy (amistre64):

area cut each time cumulative area --------------- ------------- 2(120-0+58-2) = 352, 352 2(120-2+58-4) = 344, 696 2(120-4+58-6) = 336, 1032 2(120-6+58-8) = 328, 1360 2(120-8+58-10) = 320, 1680 2(120-10+58-12) = 312, 1992 2(120-12+58-14) = 304, 2296 ....................... = 296, 2592 8 times

OpenStudy (amistre64):

8 times is really twice as much since i am counting 2 times per cycle 4 times around a complete perimeter?

OpenStudy (anonymous):

i'm saying length should be deducted by -4 after a complete cycle ....don't u think so?

OpenStudy (amistre64):

Area1 = 120*2 + 56*2 + 118*2 + 54*2 = 2(120+56+118+54) = 2(348) = 696 Area2 = 2(116+52+114+50) = 2(332) = 664 Area3 = 2(112+48+110+46) = 2(316) = 632 Total area after 3 passes around = 1992 108(2) = 216; +1992 = 2208 44(2) = 88; +2208 = 2296 42 more feet gives 84, 2296+84 = 2380 :)

OpenStudy (amistre64):

area is reduced by 32 each pass

OpenStudy (anonymous):

-2 is for one side ..hence, -4 for both sides.......as it must traverse both parallel sides of the lawn for a cycle...and 2ft is the movers length..

OpenStudy (amistre64):

ackk ... reduced by 16 i mean ;)

OpenStudy (anonymous):

ok i get it ... i was counting for only one side ....thanks amistre..:)

OpenStudy (anonymous):

so the first area should be 2(120*2 +(58-4)*2)= 348*2

OpenStudy (anonymous):

am i right?

OpenStudy (amistre64):

that is correct; you formed it differently, but its still correct :)

OpenStudy (amistre64):

now how do we determine what it means by "how many times to mow around the edges"?

OpenStudy (amistre64):

does one side = 1 pass? or do 4 sides = 1pass?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

corresponding areas 696 664 632 600 568 536....it makes 6 passes

OpenStudy (amistre64):

(600)4 = 2400 which is already too much :)

OpenStudy (amistre64):

or is that in total yard area?

OpenStudy (anonymous):

area total equals 3696 while we need 3480 so divide the last area by 4 as there are 4 sides ... we get 134 and it makes 5.5 passes..........5*4 + .5*4 = 22 corners

OpenStudy (anonymous):

mrain what do u think about the solution

OpenStudy (anonymous):

i agree thats what i got but i wasn't sure if it was correct

OpenStudy (amistre64):

...... i got my numbers for area convoluted ... :) the good thing about my age is in 3 days ill have forgotten all about it

OpenStudy (anonymous):

what grade level would you guess this question is?